OpenStudy (anonymous):

Zoie deposited $500 in her savings account at 5% compounded semiannually. WHat is her balance after 5 years?

OpenStudy (anonymous):

Total: (1+.05/2)^10=500(1.025)^10 500(1.025)^10= 640.042272